OpenStudy (anonymous):

Solve for x: 2/3 (x − 4) = 2x.

OpenStudy (anonymous):

@Sosa_ Can u plz write it using equation function......?

OpenStudy (anonymous):

\[slove for X: \frac{ 2 }{ 3} (x-4)=2x\]

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

so, now 2x*3/2=x-4 hence, 3x=x-4 so, 3x-x=-4 thus, 2x=-4 hence x=-2.......now learn to these type of linear equations by yourself.....visit khanacademy, you will understand how to solve them better and faster.

OpenStudy (dwelsher):

2/3x-8/3=2x multiply both sides by 3 2x-8=6x 4x=-8 x=-2

OpenStudy (anonymous):

@Sosa_ See the method i showed you is easier because when you bring 2/3 on the other side as 3/2, 2 on the numerator and denominator gets cancelled and you are left with 3x and then transposition becomes easier.......although it depends on what you feel easier but after a while you will be expected to mentally solve these equations, for that you may find this method easier.......